OpenStudy (anonymous):

7(e+10)=210 help me find out what e is?

OpenStudy (muhammad_nauman_umair):

e=20

OpenStudy (anonymous):

yep.

OpenStudy (muhammad_nauman_umair):

e+10=210/7 e+10=30 e=20

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

20

OpenStudy (muhammad_nauman_umair):

wlcm

OpenStudy (anonymous):

7(e+10)=210 e+10=210/7 e+10=30 e=20